In a pie chart representing a village,150degrees represent the population of children, 90degrees represents the population of men and 120degrees represent the population of women. If the children alone are 3000 In number, how many men are in the village
Answer by nerdybill(7384)

In a pie chart representing a village,150degrees represent the population of children, 90degrees represents the population of men and 120degrees represent the population of women. If the children alone are 3000 In number, how many men are in the village
.
Let x = total population of village
then
(150/360)x = 3000
(15/36)x = 3000
x = 3000(56/15)
x = 200(56)
x = 11200
.
Men in village:
(90/360)11200
(1/4)11200
2800